OpenStudy (jjuden):

Solve the equation. If necessary, round your answer to the nearest tenth. x^2 = 98

OpenStudy (anonymous):

there are two answers

OpenStudy (anonymous):

x^2-98=0 x^2 - (sqrt98)^2=0 (x-sqrt98)(x+sqrt98)=0

OpenStudy (anonymous):

x=-sqrt98 or x=-sqrt98

OpenStudy (jjuden):

49, –49 10, –10 about 9.9, –9.9 about 6.9, –6.9 these r the answers to choose from

OpenStudy (anonymous):

answer c looks closer